The surface are of a basketball is 900/pi square inches. Find the radius r of the basketball.
Found 3 solutions by Fombitz, checkley77, stanbon:
Answer by Fombitz(32388)
Answer by checkley77(12844)

SA=4PIR^
900PI=4PIR^2
R^2=900PI/4PI
R^2=225
R=SQRT225
R=15 IN.
PROOF:
900PI=4PI^15^2
900PI=4PI*225
900PI=900PI

Answer by stanbon(75887)

The surface are of a basketball is 900/pi square inches. Find the radius r of the basketball.
-----
surface area of a sphere = 4*pi*r
------
Equation:
4pi*r = 900/pi
r = 900/(4*pi^2) = 22.80 inches
=======================================
Cheers,
Stan H.
==============